# Experimental investigation of turbulence and turbulent thermal diffusion in strongly inhomogeneous and anisotropic forced convection

### v1

@@ Abstract
We investigate properties of turbulence and turbulent transport of non-inertial particles described in terms of turbulent thermal diffusion in strongly inhomogeneous and anisotropic convection forced by two turbulence generators with oscillating membrane and a steady grid in the air flow (with the Rayleigh number about $10^{8}$). Velocity field and spatial distribution of particles are measured using Particle Image Velocimetry system. The temperature distribution is measured in many locations using a temperature probe equipped with 12 E - thermocouples. In the forced convection, inhomogeneities of the mean temperature field and the particle number density in the horizontal direction in the core flow is much stronger than in the vertical direction. The mean fluid velocity structure show transition between a single-roll pattern for isothermal turbulence to double-roll patterns with increase of the temperature difference between the bottom and upper walls of the chamber. For larger temperature differences, the mean fluid velocity structure returns to a single-roll pattern. The maximum of the mean particle number density is located in the regions with minimum of the mean temperature and vise versa in agreement with the predictions based on the phenomenon of turbulent thermal diffusion.
@@ I Introduction
In an isothermal inhomogeneous turbulence, one of the important mechanism of large-scale particle clusters is turbophoresis that is combined effect of particle inertia and inhomogeneity of turbulence. This phenomenon causes formation of large-scale particle clusters in the vicinity of the minimum turbulence intensity [12, 13, 14, 15, 10, 16, 17, 18].
@@
Below we shortly discuss the mechanism of turbulent thermal diffusion of small non-inertial particles in a developed turbulence. Equation for the particle number density $n(t,{\bm{x}})$ advected by fluid velocity field ${\bm{U}}(t,{\bm{x}})$ is given by
@@
where $D=k_{B}\,T/(3\pi\rho\,\nu\,d_{\rm p})$ is the coefficient of the molecular (Brownian) diffusion of particles, $d_{\rm p}$ is the particle diameter, $\nu$ is the kinematic viscosity of the fluid, $T$ and $\rho$ are the fluid temperature and density and $k_{B}$ is the Boltzmann constant. We consider a low-Mach-number fluid flow with inhomogeneous density, so that the continuity equation in an anelastic approximation is given by $\mbox{\boldmath$\nabla$}{\bf\cdot}(\rho\,{\bm{U}})=0$.
We discuss large-scale clustering of particles. To this end, we apply a mean-field approach and use Reynolds averaging, so that all quantities are decomposed into the mean and fluctuating parts, and the fluctuating parts have zero mean values. Averaging Eq. (1) over an ensemble of turbulent velocity field, we arrive at the mean-field equation for the particle number density:
@@
where $\overline{n}=\langle n\rangle$ is the mean particle number density, $n^{\prime}$ are particle number density fluctuations, $\langle{\bm{u}}\,n^{\prime}\rangle$ is the turbulent flux of particles, $\bm{u}$} are velocity fluctuations, and the angular brackets denote ensemble averaging. For simplicity, we assume that mean velocity $\overline{\bm{U}}$ vanishes.
@@
where ${\cal Q}={\bm{\nabla}}{\bf\cdot}\left(n^{\prime}\,{\bm{u}}-\langle n^{\prime}\,{\bm{u}}\rangle\right)$ is the nonlinear term, and $({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{n}-\overline{n}({\bm{\nabla}}{\bf\cdot}{\bm{u}})$ are the source terms for particle number density fluctuations. The fluid density is decomposed into the mean fluid density $\overline{\rho}$ and fluctuations $\rho^{\prime}$, and for a low-Mach-number flow, $|\rho^{\prime}|\ll\overline{\rho}$. The anelastic approximation yields $\mbox{\boldmath$\nabla$}{\bf\cdot}{\bm{u}}\approx-\overline{\rho}^{\,-1}\,({\bm{u}}\cdot{\bm{\nabla}})\,\overline{\rho}={\bm{u}}\cdot{\bm{\lambda}}$, where ${\bm{\lambda}}=-{\bm{\nabla}}\overline{\rho}/\overline{\rho}$. Therefore, $\overline{n}({\bm{\nabla}}{\bf\cdot}{\bm{u}})=(\overline{n}/\overline{\rho}\,)({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{\rho}=-\overline{n}\,({\bm{u}}{\bf\cdot}{\bm{\lambda}})$. The ratio of the absolute values of the nonlinear term $|{\cal Q}|$ to the diffusion term $|D\Delta n^{\prime}|$ is the Péclet number for particles, that can be estimated as ${\rm Pe}=u_{0}\,\ell_{0}/D$.
@@
$\left\langle n^{\prime}\,u_{i}\right\rangle$ $=$ $-\tau_{0}\,\left\langle u_{i}u_{j}\right\rangle\,\nabla_{j}\overline{n}-\tau_{0}\,\overline{n}\,\left\langle u_{i}(\mbox{\boldmath$\nabla$}{\bf\cdot}{\bm{u}})\right\rangle$ (5)
@@
The term ${\bm{V}}^{\rm eff}\,\overline{n}$ in Eq. (5) determines the contribution to the turbulent flux of particles caused by the effective pumping velocity: ${\bm{V}}^{\rm eff}=-\tau_{0}\,\left\langle{\bm{u}}(\mbox{\boldmath$\nabla$}{\bf\cdot}{\bm{u}})\right\rangle$. Applying the anelastic approximation, ${\bm{\nabla}}{\bf\cdot}{\bm{u}}=-(1/\overline{\rho})\,({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{\rho}$, we obtain the effective pumping velocity as $V_{i}^{\rm eff}=-\tau_{0}\,\left\langle u_{i}u_{j}\right\rangle\,\lambda_{j}$, where ${\bm{\lambda}}=-{\bm{\nabla}}\overline{\rho}/\overline{\rho}$. For an isotropic turbulence, the effective pumping velocity is given by [19, 20]
@@
Various rigorous methods yield the result similar to Eq. (8), see, e.g., Ref. [7]. Equations (6)–(8) are valid also when the mean velocity does not vanish. Mean-field equation (2) for particle number density reads
where we neglect small molecular (Brownian) diffusion coefficient $D$ for large Peclet numbers. We assume that the gradients along the axis $Y$ of mean temperature, mean velocity, mean number density and turbulent diffusion coefficient are much larger than those in other directions. Thus, the steady-state solution for Eq. (9) for a zero total particle flux at the boundaries reads
Phenomenon of turbulent thermal diffusion has been studied analytically applying various theoretical methods (e.g., dimensional analysis, various $\tau$ approaches in physical and Fourier space, path integral approach, direct interaction approximation and functional approach, see [19, 20, 21, 22, 23, 24, 25, 26]). This effect has been detected in direct numerical simulations (DNS) [27, 28], observed in geophysics [29], and discussed in planetary [30] and astrophysical turbulence applications [31, 32].
Turbulent thermal diffusion has been detected in laboratory experiments with micron-size particles in temperature stratified turbulence produced by one or two oscillating grids [33, 34, 35, 26, 36, 37], or by multi-fan turbulence generator [38]. This phenomenon was even observed for nanoparticles [39] in temperature stratified turbulence. All these experiments show formation of large-scale particle clusters in the regions with the mean temperature minimum caused by turbulent thermal diffusion.
In the present paper we study in laboratory experiments properties of turbulence and a large-scale clustering of small solid particles in a convective turbulence forced by two turbulence generators with oscillating membrane and a steady grid in the air flow. This paper is organized as follows. In Sec.  II we describe experimental setup and measurement techniques. In Sec.  III we analyse the obtained experimental results, and in Sec. IV we outline conclusions.
II Experimental setup
In this section we discuss the experimental set-up and measurement technique. We study properties of turbulence and a large-scale clustering of small solid particles described in terms of the phenomenon of turbulent thermal diffusion in experiments with a convective turbulence forced by two turbulence generators with oscillating membrane and a steady grid in the air flow.
A polypropylene cone woofer pushes air through the ejector which increases the mass flow rate and velocity of the air. The outlet of the ejector is covered by a double layered Polymerized Lactic Acid square adjustable grid (24 $\times$ 24 cm2) with square array of 2-5 mm thick bars, creating a mesh structure with the mesh size $9\times 9$ mm2, which corresponds to the porosity constant $\approx 0.74$. For instance, at the frequency of 17 Hz of the oscillating membrane and r.m.s voltage of 2  V, and with a 9 mm mesh size grid, a measured turbulent velocity of up to 20 cm/s is reached at 20 mesh size distance from the steady grid.
[…]
III Experimental results
In this section we analyse the obtained experimental results in a convective turbulence forced by two turbulence generators with oscillating membrane and a steady grid attached to the side walls of the chamber (see Fig. 1). In this experiments, the turbulent kinetic energy is produced by buoyancy and the membrane oscillations. In Fig. 2 we show the mean velocity field in the core flow (excluding near-boundary regions in the chamber) for isothermal turbulence (see Fig. 2a), and forced convective turbulence at the temperature differences $\Delta T=40$ K (see Fig. 2b), $\Delta T=50$ K (see Fig. 2c) and $\Delta T=60$ K (see Fig. 2d) between the bottom and upper walls of the chamber. The mean velocity patterns in these experiments are very complicated containing one or two large-scale circulations separated by jets produced by two turbulence generators with oscillating membranes and a steady grids. In particular, we observe a transition between a single-roll pattern for isothermal turbulence (Fig. 2a) to double-roll pattern with increase of the temperature differences between the bottom and upper walls of the chamber (Figs. 2b,c), and at larger temperature differences a single-roll structure is observed again (Fig. 2d). Large-scale shear in such mean velocity fields contributes to the turbulence production rate. The imposed temperature differences $\Delta T$ between the bottom and upper walls of the chamber increases buoyancy in convective turbulence which strongly affects the mean velocity patterns. The turbulence generators in the side walls of the chamber produce jets in opposite directions creating strong shear flows.
[…]
For comparison, we also apply other approach to determine the integral turbulent length scales. In particular, we use the proper orthogonal decomposition [60, 61, 62], which allows us to extract from the measured velocity field the mean field, the part of velocity fluctuations which are strongly affected by the mean field (which characterises the coherent structures) and stochastic component of velocity fluctuations. The integral scale of turbulence is determined by a stochastic component of velocity fluctuations.
In Fig. 10 we show the distributions of the ratio $\ell_{y}^{\rm fluct}/\ell_{y}^{\rm stoch}$ of the horizontal integral turbulent length scale $\ell_{y}^{\rm fluct}$ obtained by usual Reynolds averaging (where velocity is decomposed into the mean and fluctuations) to that $\ell_{y}^{\rm stoch}$ of the stochastic part of velocity fluctuations obtained by the proper orthogonal decomposition in the core flow for isothermal turbulence and for forced convective turbulence at the various temperature differences between the bottom and upper walls of the chamber. As follows from Fig. 10 both methods yield the similar spatial distributions of the integral turbulent length scale and compared values of the ratio $\ell_{y}^{\rm fluct}/\ell_{y}^{\rm stoch}$.
In left panels of Figs. 1113, we show the distributions of the mean temperature $\overline{T}(Y,Z)$ in the core flow for forced convective turbulence at the temperature difference $\Delta T=40$ K (see Fig. 11), $\Delta T=50$ K (see Fig. 12) and $\Delta T=60$ K (see Fig. 13) between the bottom and upper walls of the chamber. We see that the inhomogeneity of the mean temperature in the horizontal $Y$ direction (along the jets) is much more stronger than in the vertical $Z$ direction.
In right panels of Figs. 1113, we show the distributions of the normalized mean particle number density $\overline{n}(Y,Z)/\overline{n}_{0}$. As expected, the inhomogeneity of the particle number density in the horizontal direction is much more stronger than in the vertical direction, and the maximum of the mean particle number density is located in the regions with minimum of the mean temperature $\overline{T}(Y,Z)$ in accordance with the phenomenon of turbulent thermal diffusion (see Figs. 1113).
In Fig. 14 we plot the dependencies of the mean temperature $\overline{T}(Y)/\overline{T}_{0}$ and the normalized mean particle number density $\overline{n}(Y)/\overline{n}_{0}$ on the horizontal coordinate $Y$ in the core flow for three fixed vertical positions: $Z/L_{z}=0.23$ (diamonds), $Z/L_{z}=0.4$ (circles), $Z/L_{z}=0.58$ (crosses), for forced convective turbulence at different temperature differences between the bottom and upper walls of the chamber. In spite of the presence of the large-scale circulations superimposed to the convective turbulence, Fig. 14 demonstrates the tendency of the localization of the maximum of the mean particle number density in the vicinity of the minimum in the mean temperature distribution, and vise versa.
Note that the terminal fall velocity for the micron-size particles is about $10^{-2}$ cm/s, while the turbulent velocity in the experiments with the forced convective turbulence is much larger than the particle terminal fall velocity (it is about $25$ cm/s near the turbulence generator and is more than $10$ cm/s far from the turbulence generator). Also the mean fluid velocity in several times smaller than the turbulent velocity. This implies that sedimentation of particles does not play any role in a formation of inhomogeneous particle distributions near the bottom wall of the chamber. The effective pumping velocity caused by turbulent thermal diffusion is about 1155 cm/s. This implies that the turbulent and effective pumping velocities in the experiments are much larger than the terminal fall velocity for micron-size particles.
IV Conclusions
Properties of turbulence and turbulent transport of non-inertial particles have been studied in experiments in a turbulent convection forced by two turbulence generators at the side walls of the chamber with oscillating membrane and a steady grid in the air flow with the Rayleigh number about $10^{8}$. Particle Image Velocimetry system has been used for measurements of velocity fields and mean particle number density spatial distributions, while the temperature field has been measured in many locations applying a temperature probe equipped with 12 E - thermocouples.
These experiments show that distributions of the mean particle number density are anti-correlated with the mean temperature distributions. In particular, the maximum particle number density is located in the regions with the minimum mean fluid temperature, which is qualitatively consistent with the predictions based on the phenomenon of turbulent thermal diffusion. Deviations from this trend occurs in regions with strong vertical mean fluid motions.
AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose.
DATA AVAILABILITY

### v2

@@ Abstract
We investigate properties of turbulence and turbulent transport of non-inertial particles described in terms of turbulent thermal diffusion in strongly inhomogeneous and anisotropic convection forced by two similar turbulence generators with oscillating membrane and a steady grid in the air flow (with the Rayleigh number about $10^{8}$). Velocity field and spatial distribution of particles are measured using Particle Image Velocimetry system. The temperature distribution is measured in many locations using a temperature probe equipped with 12 E - thermocouples. In the forced convection, the gradients of the mean temperature field and the particle number density in the horizontal direction in the core flow are much stronger than in the vertical direction. The mean fluid velocity structure show transition between a single-roll pattern for isothermal turbulence to double-roll patterns with increase of the temperature difference between the bottom and upper walls of the chamber. For larger temperature differences, the mean fluid velocity structure returns to a single-roll pattern. In the turbulent regions with large mean temperature gradients, the dominant effect of the large-scale particle clustering is turbulent thermal diffusion, resulting in that the maximum of the mean particle number density is located in the regions with minimum of the mean temperature and vise versa. Deviations from this feature is observed in the regions with strong mean fluid velocities where the mean temperature gradients are small.
@@ I Introduction
In an isothermal inhomogeneous turbulence, one of the important mechanism of large-scale particle clusters is turbophoresis that is combined effect of particle inertia and inhomogeneity of turbulence. This phenomenon causes the formation of large-scale particle clusters in the vicinity of the minimum turbulence intensity [12, 13, 14, 15, 10, 16, 17, 18].
@@
Below we shortly discuss the mechanism of turbulent thermal diffusion resulting in the large-scale clustering of small non-inertial particles in a developed turbulence. Equation for the particle number density $n(t,{\bm{x}})$ advected by fluid velocity field ${\bm{U}}(t,{\bm{x}})$ is given by
@@
where $D=k_{B}\,T/(3\pi\rho\,\nu\,d_{\rm p})$ is the coefficient of the molecular (Brownian) diffusion of particles, $d_{\rm p}$ is the particle diameter, $\nu$ is the kinematic viscosity of the fluid, $T$ and $\rho$ are the fluid temperature and density and $k_{B}$ is the Boltzmann constant. We consider a low-Mach-number fluid flow with inhomogeneous density, so that the continuity equation in an anelastic approximation is given by $\mbox{$\nabla$}{\bf\cdot}(\rho\,{\bm{U}})=0$.
To describe the large-scale clustering of particles, we apply a mean-field approach and use Reynolds averaging, so that all quantities are decomposed into the mean and fluctuating parts, and the fluctuating parts have zero mean values. Averaging Eq. (1) over an ensemble of turbulent velocity field, we arrive at the mean-field equation for the particle number density:
@@
where $\overline{n}=\langle n\rangle$ is the mean particle number density, $n^{\prime}$ are particle number density fluctuations, $\langle{\bm{u}}\,n^{\prime}\rangle$ is the turbulent flux of particles, $\bm{u}$} are velocity fluctuations, and the angular brackets denote ensemble averaging. The particle turbulent flux $\langle{\bm{u}}\,n^{\prime}\rangle$ is given by [19, 20]
$\left\langle{\bm{u}}\,n^{\prime}\right\rangle={\bm{V}}^{\rm eff}\,\overline{n}-D_{\rm T}\,{\bm{\nabla}}\overline{n},$ (3)
where $D_{\rm T}$ is the turbulent diffusion coefficient and ${\bm{V}}^{\rm eff}$ is the effective pumping velocity that for small non-inertial particles is given by
${\bm{V}}^{\rm eff}=D_{\rm T}\,{{\bm{\nabla}}\overline{\rho}\over\overline{\rho}}\approx-D_{\rm T}\,{{\bm{\nabla}}\overline{T}\over\overline{T}}.$ (4)
Here $\overline{\rho}$ and $\overline{T}$ are the mean fluid density and temperature, respectively, and we take into account that for a small mean pressure gradient, ${\bm{\nabla}}\,\ln\overline{\rho}\approx-{\bm{\nabla}}\ln\overline{T}$. Note that various rigorous methods yield Eqs. (3)–(4), which implies that this result is robust (see, e.g., Ref. [7]). For illustration, the equation for the turbulent flux of particles $\langle{\bm{u}}\,n^{\prime}\rangle$ for an isotropic turbulence is derived in Appendix. Using Eqs. (3)–(4), we rewrite Eq. (2) for the mean particle number density as
where we neglect small molecular (Brownian) diffusion coefficient $D$ for large Peclet numbers. Note that Eqs. (3)–(4) are also valid when the mean velocity does not vanish. We assume that the gradients along the horizontal axis $Y$ of the mean temperature, mean velocity, mean number density and turbulent diffusion coefficient are much larger than those in other directions. This condition corresponds to that observed in the experiments discussed in this paper. Thus, the steady-state solution for Eq. (5) for a zero total particle flux at the boundaries reads
Phenomenon of turbulent thermal diffusion has been studied analytically applying various theoretical methods (e.g., the dimensional analysis, the $\tau$ approaches applied in physical and Fourier spaces, the path integral approach, the direct interaction approximation and functional approach, see [7, 19, 20, 21, 22, 23, 24, 25, 26]). This effect has been detected in direct numerical simulations (DNS) [27, 28], observed in geophysics [29], and discussed in planetary [30] and astrophysical turbulence applications [31, 32].
Turbulent thermal diffusion has been detected in laboratory experiments with micron-size particles in temperature stratified turbulence produced by one or two oscillating grids [33, 34, 35, 26, 36, 37], or by multi-fan turbulence generator [38]. This phenomenon was even observed for nanoparticles [39] in temperature stratified turbulence. All these experiments show the formation of large-scale particle clusters in the regions with the mean temperature minimum caused by turbulent thermal diffusion.
In the present paper we study in laboratory experiments properties of turbulence and a large-scale clustering of small solid particles in a convective turbulence forced by two similar turbulence generators which contain both, oscillating membrane and a steady grid in the air flow. This forcing is completely different from that applied to turbulent convection in the previous experiments [34, 35, 37], where the large-scale circulations in turbulent convection were destroyed by the applied forcing, so that the mean velocity fields were much smaller than velocity fluctuations. In the present convective experiments, the forcing produces very complicated large-scale velocity field with large-scale circulations and strong horizontal jets. As the result, the gradients of the mean temperature and the mean particle number density in the horizontal direction occur to be much larger than in the vertical direction. Thus, in the first time, the large-scale clustering of small solid particles is investigated in the horizontal direction rather than in the vertical direction.
This paper is organized as follows. In Sec.  II we describe experimental setup and measurement techniques. In Sec.  III we analyse the obtained experimental results, and in Sec. IV we outline conclusions. In Appendix we derive the equation for the turbulent flux of particles for an isotropic turbulence.
II Experimental setup
In this section we discuss the experimental set-up and measurement technique. We study properties of turbulence and a large-scale clustering of small solid particles described in terms of the phenomenon of turbulent thermal diffusion in experiments with a convective turbulence forced by two similar turbulence generators with oscillating membrane and a steady grid in the air flow.
A polypropylene cone woofer pushes air through the ejector which increases the mass flow rate and velocity of the air. The outlet of the ejector is covered by a double layered Polymerized Lactic Acid square adjustable grid (24 $\times$ 24 cm2) with square array of 2-5 mm thick bars, creating a mesh structure with the mesh size $9\times 9$ mm2, which corresponds to the porosity constant $\approx 0.74$ (see Fig. 1). For instance, at the frequency of 17 Hz of the oscillating membrane and r.m.s voltage of 2  V, and with a 9 mm mesh size grid, a measured turbulent velocity of up to 20 cm/s is reached at 20 mesh size distance from the steady grid.
[…]
III Experimental results
In this section we analyse the obtained experimental results in a convective turbulence forced by two similar turbulence generators with oscillating membrane and a steady grid attached to the side walls of the chamber (see Fig. 1). In Fig. 2 we show the mean velocity field in the core flow (excluding near-boundary regions in the chamber) for isothermal turbulence (see Fig. 2a), and forced convective turbulence at the temperature differences $\Delta T=40$ K (see Fig. 2b), $\Delta T=50$ K (see Fig. 2c) and $\Delta T=60$ K (see Fig. 2d) between the bottom and upper walls of the chamber. The mean velocity patterns in these experiments are very complicated containing one or two large-scale circulations separated by jets produced by two turbulence generators with oscillating membranes and a steady grids. In particular, we observe a transition between a single-roll pattern for isothermal turbulence (Fig. 2a) to double-roll pattern with increase of the temperature differences between the bottom and upper walls of the chamber (Figs. 2b,c), and at larger temperature differences a single-roll structure is observed again (Fig. 2d).
[…]
For comparison, we also apply other approach to determine the integral turbulent length scales. In particular, we use the proper orthogonal decomposition [53, 54, 55], which allows us to extract from the measured velocity field the mean field, the part of velocity fluctuations which are strongly affected by the mean field (which characterises the coherent structures) and stochastic component of velocity fluctuations. The integral scale of turbulence is determined by a stochastic component of velocity fluctuations. In Fig. 9 we show the distributions of the ratio $\ell_{y}^{\rm fluct}/\ell_{y}^{\rm stoch}$ of the horizontal integral turbulent length scale $\ell_{y}^{\rm fluct}$ obtained by usual Reynolds averaging (where velocity is decomposed into the mean and fluctuations) to that $\ell_{y}^{\rm stoch}$ of the stochastic part of velocity fluctuations obtained by the proper orthogonal decomposition in the core flow for isothermal turbulence and for forced convective turbulence at the various temperature differences between the bottom and upper walls of the chamber.
As follows from Fig. 9, in many regions both methods yield the similar spatial distributions of the integral turbulent length scale and compared values of the ratio $\ell_{y}^{\rm fluct}/\ell_{y}^{\rm stoch}$. Only there are some spatial regions, where this ratio exceeds 1 substantially. May be this anomaly is related to the fact that in these regions there is a very complicated mean velocity field and the proper orthogonal decomposition does not work well there.
In left panels of Figs. 1113, we show the distributions of the mean temperature $\overline{T}(Y,Z)$ in the core flow for forced convective turbulence at the temperature difference $\Delta T=40$ K (see Fig. 11), $\Delta T=50$ K (see Fig. 12) and $\Delta T=60$ K (see Fig. 13) between the bottom and upper walls of the chamber. We see that the gradients of the mean temperature in the horizontal $Y$ direction (along the jets) are much more stronger than in the vertical $Z$ direction.
In right panels of Figs. 1113, we show the distributions of the normalized mean particle number density $\overline{n}(Y,Z)/\overline{n}_{0}$. As expected, the gradients of the particle number density in the horizontal direction are much more stronger than in the vertical direction. In the regions with the strong mean temperature gradients, the maximum of the mean particle number density is located in the regions with minimum of the mean temperature $\overline{T}(Y,Z)$ in accordance with the phenomenon of turbulent thermal diffusion (see Figs. 1113).
In Fig. 14 we plot the dependencies of the mean temperature $\overline{T}(Y)/\overline{T}_{0}$ and the normalized mean particle number density $\overline{n}(Y)/\overline{n}_{0}$ on the horizontal coordinate $Y$ in the core flow for three fixed vertical positions: $Z/L_{z}=0.23$ (diamonds), $Z/L_{z}=0.4$ (circles), $Z/L_{z}=0.58$ (crosses), for forced convective turbulence at different temperature differences between the bottom and upper walls of the chamber. In spite of the presence of the large-scale circulations superimposed to the convective turbulence, Fig. 14 demonstrates the tendency of the localization of the maximum of the mean particle number density in the vicinity of the minimum in the mean temperature distribution, and vise versa. However, this conclusion is only valid in the turbulent regions with large mean temperature gradients. For these regions, the dominant effect of large-scale particle clustering is turbulent thermal diffusion. Deviations from this trend occurs in the regions with strong mean fluid motions where the mean temperature gradient is small.
Note that the terminal fall velocity for the micron-size particles is about $10^{-2}$ cm/s, while the turbulent velocity in the experiments with the forced convective turbulence is much larger than the particle terminal fall velocity (it is about $25$ cm/s near the turbulence generator and is more than $10$ cm/s far from the turbulence generator). This implies that sedimentation of particles does not play any role in a formation of inhomogeneous particle distributions. The effective pumping velocity caused by turbulent thermal diffusion is about 1155 cm/s. This implies that the turbulent and effective pumping velocities in the experiments are much larger than the terminal fall velocity for micron-size particles.
IV Conclusions
Properties of turbulence and turbulent transport of non-inertial particles have been studied in experiments in a turbulent convection forced by two similar turbulence generators at the side walls of the chamber with oscillating membrane and a steady grid in the air flow with the Rayleigh number about $10^{8}$. Particle Image Velocimetry system has been used for measurements of velocity fields and mean particle number density spatial distributions, while the temperature field has been measured in many locations applying a temperature probe equipped with 12 E - thermocouples.
In the turbulent regions with large mean temperature gradient, the dominant effect of large-scale particle clustering is turbulent thermal diffusion so that the distributions of the mean particle number density are anti-correlated with the mean temperature distributions. This implies that the maximum particle number density is located in the regions with the minimum mean fluid temperature, which is qualitatively consistent with the predictions based on the phenomenon of turbulent thermal diffusion. However, deviations from this trend occurs in the regions with strong mean fluid motions where the mean temperature gradient is small.
Note that we have performed experiments with small non-inertial particles. For inertial particles a large-scale clustering can be more stronger, because there is an additional positive contribution to the effective pumping velocity (see, e.g., Ref. [7]). However, it is very difficult to perform experiments with inertial particles that is a subject of a separate investigation.
In the present study we have done only a qualitative comparison between theoretical predictions of large-scale particle clustering and experimental results. Since we measure a two-dimensional velocity field (in the $YZ$ plane), and since the large-scale complicated three-dimensional flow contributes to the large-scale clustering of particles in the turbulent regions with small mean temperature gradients, it is not visible to perform a quantitative comparisons between theoretical and experimental results.
ACKNOWLEDGEMENTS
We are thankful to the two referees for providing constructive comments that improved our paper.
AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose.
DATA AVAILABILITY
Appendix A Derivation of the turbulent flux of particles
@@
where ${\cal Q}={\bm{\nabla}}{\bf\cdot}\left(n^{\prime}\,{\bm{u}}-\langle n^{\prime}\,{\bm{u}}\rangle\right)$ is the nonlinear term, and $({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{n}-\overline{n}({\bm{\nabla}}{\bf\cdot}{\bm{u}})$ are the source terms for particle number density fluctuations. Here, for simplicity, we assumed that mean velocity $\overline{\bm{U}}$ vanishes. The fluid density is decomposed into the mean fluid density $\overline{\rho}$ and fluctuations $\rho^{\prime}$, and for a low-Mach-number flow, $|\rho^{\prime}|\ll\overline{\rho}$. The anelastic approximation yields $\mbox{$\nabla$}{\bf\cdot}{\bm{u}}\approx-\overline{\rho}^{\,-1}\,({\bm{u}}\cdot{\bm{\nabla}})\,\overline{\rho}={\bm{u}}\cdot{\bm{\lambda}}$, where ${\bm{\lambda}}=-{\bm{\nabla}}\overline{\rho}/\overline{\rho}$. Therefore, $\overline{n}({\bm{\nabla}}{\bf\cdot}{\bm{u}})=(\overline{n}/\overline{\rho}\,)({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{\rho}=-\overline{n}\,({\bm{u}}{\bf\cdot}{\bm{\lambda}})$. The ratio of the absolute values of the nonlinear term $|{\cal Q}|$ to the diffusion term $|D\Delta n^{\prime}|$ is the Péclet number for particles, that can be estimated as ${\rm Pe}=u_{0}\,\ell_{0}/D$.
@@
$\left\langle n^{\prime}\,u_{i}\right\rangle$ $=$ $-\tau_{0}\,\left\langle u_{i}u_{j}\right\rangle\,\nabla_{j}\overline{n}-\tau_{0}\,\overline{n}\,\left\langle u_{i}(\mbox{$\nabla$}{\bf\cdot}{\bm{u}})\right\rangle$ (9)
@@
The term ${\bm{V}}^{\rm eff}\,\overline{n}$ in Eq. (9) determines the contribution to the turbulent flux of particles caused by the effective pumping velocity: ${\bm{V}}^{\rm eff}=-\tau_{0}\,\left\langle{\bm{u}}(\mbox{$\nabla$}{\bf\cdot}{\bm{u}})\right\rangle$. Applying the anelastic approximation, ${\bm{\nabla}}{\bf\cdot}{\bm{u}}=-(1/\overline{\rho})\,({\bm{u}}{\bf\cdot}{\bm{\nabla}})\overline{\rho}$, we obtain the effective pumping velocity as $V_{i}^{\rm eff}=-\tau_{0}\,\left\langle u_{i}u_{j}\right\rangle\,\lambda_{j}$, where ${\bm{\lambda}}=-{\bm{\nabla}}\overline{\rho}/\overline{\rho}$. For an isotropic turbulence, the effective pumping velocity is given by [19, 20]